### v1

# Adaptive Group Policy Optimization: Towards Stable Training and Token-Efficient Reasoning

Abstract
Since DeepSeek-R1 popularized, Group Relative Policy Optimization (GRPO) has become the core part of Reasoning LLMs training. However, we find some deficiency that influences RL stability and inference efficiency. Thus, we propose Adaptive Group Policy Optimization (AGPO) which contains two simple but effective modifications: a revised advantage estimation method to mitigate zero-variance situations; a length-based reward, incentivizing the model to avoid overthinking. The experiments demonstrate our methods achieve more stable training and comparable or superior performance with significantly fewer tokens in reasoning steps.
This is an unfinished version and will be updated. We aim to share some findings.
Adaptive Group Policy Optimization: Towards Stable Training and Token-Efficient Reasoning
Chen Li, Nazhou Liu, Kai Yang HPC-AI Tech {lichen,liunazhou,yangkai}@luchentech.com
1 Introduction
Large Language Models (LLMs)Bommasani et al. (2021); Wei et al. (2022); Zhao et al. (2023) have achieved impressive performance through extensive pre-training and fine-tuning processes. However, effectively generating desired model responses often necessitates aligning outputs with specific downstream tasks and human preferencesWang et al. (2023); Wolf et al. (2023).
For alignment challenges, reinforcement learning from human feedback (RLHF)Bai et al. (2022); Kaufmann et al. (2023) is introduced as a prominent post-training strategy, adopted by notable LLMs including GPT-4, Claude, Gemini, and DeepSeek. They have explored various optimization techniques such as Proximal Policy Optimization (PPO)Schulman et al. (2017). Recently, to significantly reduce computational and memory overhead associated with PPO, DeepSeek eliminated the value model and proposed Group Relative Policy Optimization (GRPO)Guo et al. (2025), which achieves high computational efficiency and excellent reasoning performance, surpassing other open-source models ranging from 7B to 70B.
Despite the demonstrated success of GRPO, it introduces challenges that can affect stable training and inference efficiency.
(1) Vanished Advantage Signal: GRPO computes the normalized advantage $A_{i}=\frac{r_{i}-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}}$ for each sampled output $i$ where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of group rewards. However, when all rewards within a group are identical, this normalized numerator and denominator approach 0, causing the advantage signal to vanish, which potentially stalling training.
(2) Inefficient CoT Length: While lengthy As GRPO lacks mechanisms to discourage excessively long chain-of-thought (CoT), models tend to produce overly verbose explanations. A refined approach that rewards concise, effective reasoning is essential to improve token efficiency.
To address these issues, we propose an enhanced GRPO training algorithm, Adaptive Group Policy Optimization (AGPO). Our main contributions are summarized as follows:
  • Training Efficiency: By identifying the limitations of GRPO’s advantage, we introduce a modified advantage estimation method that addresses zero-variance scenarios, ensuring continuous and effective learning.
  • Token Efficiency: We incorporate a self-adaptive length reward into RL training. Compared with GRPO baselines, our approach achieves comparable performance with significantly fewer response tokens.
2 Background
2.1 Policy Gradient
Policy gradient method is one of the most fundamental RL algorithm which directly model and optimize the policy. For any differentiable policy, the policy gradient is :
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})A_{t}\right]$ (1)
Where advantage $A_{t}$ is the most crucial part for policy gradient method, which defines how much better a specific action $a_{t}$ is compared to average action give a state $s_{t}$.
2.2 Proximal Policy Optimization (PPO)
PPO (Schulman et al., 2017) is one of policy gradient methods which uses clipped surrogate objective for policy optimization. Specifically, it maximize the following objective:
$J_{\text{PPO}}(\theta)=\mathbb{E}_{\pi_{\theta_{\text{old}}}}\left[\min\left(r_{t}{A}_{t},\operatorname{clip}(r_{t},1-\epsilon,1+\epsilon){A}_{t}\right)\right]$ (2)
Where $\epsilon$ is a hyper-parameter used for tuning clipping range. $A_{t}$ is the advantage, which generally will be computed by utilizing Generalized Advantage Estimation (GAE) (Schulman et al., 2015) in PPO. $r_{t}$ is the probability ratio of predicting token $o_{t}$ for a given question $q$ before and after the policy update:
$r_{t}(\theta)=\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}$ (3)
2.3 Group Relative Policy Optimization (GRPO)
Compared to PPO, GRPO (Shao et al., 2024) significantly saves the training cost through eliminating the critic model in PPO. This is achieved by approximating the advantage $A_{i}$ as group-normalized reward:
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})}$ (4)
Where $\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})$ and $\text{std}(\{r_{1},r_{2},\dots,r_{G}\})$ denotes the within-group mean and standard deviation respectively.
With the estimated advantage and a KL divergence penalty term, GRPO generates a group of outputs $\{o_{i}\}_{i=1}^{G}$ based on $\pi_{\theta_{\text{old}}}$ for each question $q$ and update $\pi_{\theta}$ with following objective:
$J_{\text{GRPO}}(\theta)&=\mathbb{E}_{(q)\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}{A}_{i},\\ &\quad\quad\quad\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\right){A}_{i}\Bigg{)}-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Bigg{)}\Bigg{]}\end{aligned}$
Where $\epsilon$ and $\beta$ are hyper-parameters.
3 Adaptive Group Policy Optimization
During RL training with the GRPO algorithm, we find two useful modifications to improve efficiency and performance.
3.1 Modification 1: Corner case of Advantage estimation
It is easy to find that if the rewards of the group are all equal, like all correct or all wrong, then the advantages become all zero. Therefore, the corresponding sample in the batch has no effect in the training. As model training progresses, this situation increases. We believe that it is beneficial to incorporate the group accuracy signal into the training. The new advantage function is as follows:
$A_{i}=\left\{\begin{array}[]{ll}1,&r_{mean}=r_{max}\\ -1,&r_{mean}=r_{min}\\ \frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})},&\text{otherwise}\end{array}\right.$ (6)
By replacing the original GRPO with Equation 6, the objective maintains all information in the batch, including the directions that policy models follow to reason better.
3.2 Modification 2: Length Reward
It is noticeable that reasoning models often have overthinking problem (Kumar et al., 2025). To mitigate it and improve token efficiency, we applied the length reward described in (Team et al., 2025) by adding it to the accuracy reward, but the performance got severe decline during the training. Hence, common length rewards are not suitable for GRPO settings. We design a simple but effective method that dramatically reduces response length while achieving same model effect. The length reward is defined as follows:
$r_{len}(i)=\left\{\begin{array}[]{ll}0,&\text{if }\text{mean}(r_{acc})=r_{max}\\ \lambda,&\text{otherwise}\end{array}\right.$ (7)
where,
$\lambda=\left\{\begin{array}[]{ll}0,&r_{acc}(i)=r_{min}\\ 1-\frac{len(o_{i})-min_{i}len(o_{i})}{max_{i}len(o_{i})-min_{i}len(o_{i})},&r_{acc}(i)=r_{max}\end{array}\right.$
In Equation 7, we set length reward as 0 when outputs are all correct to avoid negative weighted mean advantages. We only use $i\in correct\_set{\{o_{1},o_{2},\cdots,o_{G}\}}$ when calculate $max_{i}len(o_{i})$ and $min_{i}len(o_{i})$ in Equation 8. The final reward is $r=r_{acc}+\gamma r_{len}$.
4 Experiment
We conduct a few experiments for evaluating how these modifications affect the RL training of Reasoning models.
4.1 Implementation Details
We use Qwen2.5-Math-7B as the base model. All experiments are conducted on the MATH train set (Hendrycks et al., 2021) and tested on MATH500 (Lightman et al., 2023) with the metric of Pass@1. VeRL (Sheng et al., 2024) is utilized to perform RL training on the base model with a mini batch size of 32 and a learning rate of $61e-$ 6. The number of group rollout is 5. As for reward settings, the $r_{min}$ and $r_{max}$ for accuracy reward are set to 0 and 1 respectively. $\gamma$, the coefficient of length reward, is $0.1$.
4.2 Training Efficiency
Figure 1: Policy loss curves of GRPO and GRPO m1(modification 1)
Figure 2: Test set performance of GRPO and GRPO m1
After applying modification 1, we observe a more consistent and continuously decreased loss compared to the original GRPO from Figure 1. It provides the training more sufficient learning signals from batch data. Figure 2 also shows that the performance of GRPO-m1 is more stable at the early stage and exceeds the performance of GRPO most of the time. They achieve 78.2 and 76.2 at the best checkpoint respectively.
4.3 Token Efficiency
Model Overall Level 1 Level 2 Level 3 Level 4 Level 5
Qwen2.5-Math-7B 61.0 (620) 86.0 (299) 82.2 (454) 67.6 (584) 57.0 (632) 37.3 (853)
Qwen2.5-Math-7B-GRPO 77.2 (640) 97.7 (354) 86.7 (437) 89.5 (521) 76.4 (707) 55.7 (863)
Qwen2.5-Math-7B-AGPO 77.2 (463) 97.7 (231) 88.9 (327) 91.4 (369) 72.4 (507) 55.7 (671)
Table 1: Comparison of performance and token consumption on MATH500. We record results of overall and different difficulty levels with the format of pass@1 (average response tokens).
We add modification 2 on top of GRPO-m1 in order to further improve token efficiency. As Table 1 shows, our method dramatically reduces response length from 640 to 463 while keeping on-par performance. There is a $22\%$-$35\%$ improvement in token efficiency among difficulty levels. For validating fine-grained results of difficulty levels, we use evaluation codes of OpenR (Wang et al., 2024) after training finished, so the accuracy values may have a slight difference than those in subsection 4.2.
Limitations
We will experiment on different base models and datasets in future to validate universality of our methods. More ablation studies around modifications will be taken as well. Besides, we plan to generalize modification 1 to all groups in a batch and adjust pre-defined advantage values of corner cases.